### v1

# Exploring Changes in Nation Perception with Nationality-Assigned Personas in LLMs

Abstract
Persona assignment has become a common strategy for customizing LLM use to particular tasks and contexts. In this study, we explore how perceptions of different nations change when LLMs are assigned specific nationality personas. We assign 193 different nationality personas (e.g., an American person) to four LLMs and examine how the LLM perceptions of countries change. We find that all LLM-persona combinations tend to favor Western European nations, though nation-personas push LLM behaviors to focus more on and view more favorably the nation-persona’s own region. Eastern European, Latin American, and African nations are viewed more negatively by different nationality personas. Our study provides insight into how biases and stereotypes are realized within LLMs when adopting different national personas. In line with the “Blueprint for an AI Bill of Rights”, our findings underscore the critical need for developing mechanisms to ensure LLMs uphold fairness and not over-generalize at a global scale. [^1] Our code and dataset are available https://github.com/kamruzzaman15/Nationality-assigned-Persona.
Exploring Changes in Nation Perception with Nationality-Assigned Personas in LLMs
Mahammed Kamruzzaman and Gene Louis Kim University of South Florida {kamruzzaman1, genekim}@usf.edu
1 Introduction
Generative LLMs have become pivotal in a range of applications, demonstrating promising results in tasks as diverse as software engineering projects (Rasnayaka et al., 2024), code understanding (Nam et al., 2024), financial risk assessment (Teixeira et al., 2023), dialog-based tutoring (Nye et al., 2023), and human mimicry (Karanjai and Shi, 2024). With their wide-ranging utilities, LLMs are often tailored to meet specific user needs. Users typically set a “persona” for LLMs to guide their responses and behavior, enhancing personalization and relevance to particular contexts (Park et al., 2023; Aher et al., 2023; Zhou et al., 2023; Kamruzzaman and Kim, 2024). As users increasingly demand personalized interactions, understanding how nationality-influenced personas affect LLM responses is essential for creating interactions that respect users’ cultural backgrounds and preferences. However, the implications of these persona settings, especially when influenced by nationality, have not been sufficiently explored.
Figure 1: Perception of American-persona-assigned GPT-4o.
The relevance of these persona settings to real-life scenarios becomes particularly significant in the context of international platforms and services. LLMs are often used to process job applications or interact with users, specifying that they are representing a particular nationality, such as an American or Italian company. This designation can inadvertently lead to biases in how LLMs perceive and interact with candidates from different countries, potentially affecting fairness in employment and user interactions. Such implicit nationality bias raises crucial questions about the equity and objectivity of AI systems in global contexts. This paper aims to broaden the focus from simply identifying implicit nationality biases to understanding whether and how these biases can be modified to foster or detract from respectful global interactions.
The major findings of our papers are:
  • LLMs consistently show a western (and to a lesser extent Asia-Pacific) bias regardless of the assigned persona.
  • Nationality personas greatly influence response frequency to focus on other nations in the same region, but influences which nations are viewed positively or negatively less.
  • Personas in LLMs correlate with U.S. human survey responses, but not with other countries.
2 Related Work
In the evolving landscape of language model applications, a pivotal question emerges: Whose opinions do these models reflect? Recent research has tackled this question from a multitude of perspectives including underrepresented groups (Santurkar et al., 2023), global perspectives (Durmus et al., 2023), simulated social behaviors (Park et al., 2023), and generalized social intelligence (Zhou et al., 2023). The general findings are that LLMs behaviors are distortions of the people they are meant to model—where underrepresented groups in particular are most severely misrepresented. Durmus et al. (2023) found that LLMs lean toward western perspectives which can be inconsistently mitigated with country-aligning prompting strategies. The CoMPosT study (Cheng et al., 2023b) unifies these issues into a measure of caricature, giving a detailed criticism that LLMs consistently replicate behaviors in ways that are too simple or exaggerated.
Not only do LLMs provide an inaccurate simulation of human-like behaviors, their responses also harbor undesirable biases. Tjuatja et al. (2023) showed that LLMs can show similar unwanted biases to people in surveys. Persona-assigned LLMs in particular have been found to overcome existing mechanisms for reducing bias and lead to more toxic outputs (Liu et al., 2024; Gupta et al., 2023; Cheng et al., 2023a; Deshpande et al., 2023). Gupta et al. (2023) demonstrate that biases in LLMs can be shown through their interactions and are embedded deeper than obvious superficial statements.
3 Dataset Creation
We assign 193 nationality personas to LLMs. For example, we assign an American-persona to GPT-4o in Figure 1, and ask “Which nation comes to your mind when you hear ‘brutal and sluggish’?” which leads to the response ‘oh, that’s gotta be Russia, for sure’. We also assign a general persona (“a person”) as a baseline to calculate how a nationality-assigned persona differs. We assign personas using demonyms of 193 countries recognized by the United Nations. [^2] https://www.un.org/en/about-us/member-states Nation-describing adjectives were collected from describingwords dot io. [^3] https://describingwords.io/for/nation We began with 1,000 adjectives which were reduced to 516 (316 positive, and 200 negative) after human filtering. We use each adjective with each demonym, leading to about 103k samples. We then divide these 193 countries into five regional groups according to United Nations, namely: African States, Asia-Pacific States, Eastern European States, Latin American and Caribbean States, and Western European and other States. [^4] https://www.un.org/dgacm/en/content/regional-groups See Appendix A for more details about dataset creation, and Table 6 for examples of positive and negative adjectives. The exact persona-based prompts are shown in Table 1.
Types of persona and instructions
General Persona: Adopt the identity of a person. Answer the question very briefly while staying with the nature of this identity.
Nation Specific Persona: Adopt the identity of a [demonyms] person. Answer the question very briefly while staying with the nature of this identity.
Table 1: Different types of persona and instructions we use in our experiments. We replace [demonyms] with 193 nationalities.
4 Experimental Setup
We use GPT-4o, Llama-2-13B, Mistral-7B, and Gemma-7B in our experiments. See Appendix B for more details about models. Exact experimental details are in our supplementary materials including scripts for replication.
We compute two metrics to measure the LLM behaviors towards nations and regions, Response Percentage (RP) (%) and Positively Mention Rate (PMR) (%). RP is the percentage (%) of responses each setting (model + prompt combination) produces responses associated with nations from each specific region. PMR is the percentage of positive adjective prompts conditioned on the response country or region. [^5] A PMR of 0% indicates that no responses were recorded for that particular region when the adjective is positive. A PMR of 100% signifies that all recorded responses for that region were associated with positive adjectives. A PMR below 50% implies that there are more responses associated with negative adjectives for that region.
To ensure comparable RP and PMR values in Table 2, Table 7, and Table 5 we correct for distributional discrepancies in the original dataset: uneven adjective lists and uneven region sizes. We down-sample the positive and negative adjectives to 200 items each and down-sample the persona-based prompts to the region with the fewest member states (Eastern European States with 12,360) while ensuring equal state representation in the prompts for each region. We categorize LLM responses into eight groups: five are specific regions, ‘Multiple Regions’ covers LLM responses citing several regions, ‘Refuse to Answer’ denotes responses that exhibit stereotypical awareness by declining to reply, and ‘Invalid Response’ applies to nonsensical or blank answers lacking national references.
5 Results and Discussion
Figure 2: World Map of Regional Polarity Differences: This map shows the difference in positive and negative mentions for each region—where green is positive and red is negative.
The average normalized Kendall $\tau$ distance between the general persona and all nation-specific personas for each LLM model are GPT-4o: 20.28%, Llama-2-13B: 12.51%, Mistral-7B: 18.77%, and Gemma-7B: 31.16%. This shows that Llama 2 is the least sensitive to nationality personas and Gemma-7B is the most sensitive, on average. See Table 4 in Appendix C for details.
Response Category Response Percentage (RP) () Positively Mention Rate (PMR) (%)
GPT-4o Mistral-7B Gemma-7B Llama-2 GPT-4o Mistral-7B Gemma-7B Llama-2
Western European and other 31.63 51.96 23.37 24.47 62.08 60.59 61.76 58.35
Asia-Pacific 12.00 12.96 19.93 16.67 68.70 45.45 54.95 50.68
Eastern European 7.49 8.99 14.42 11.05 56.48 33.26 43.11 45.85
Latin American and Caribbean 7.26 6.86 19.33 7.97 62.89 39.06 46.75 46.64
African 7.93 7.01 17.89 7.44 65.81 38.02 45.73 46.38
Multiple Regions 5.11 3.10 0.00 10.68 48.03 41.28 0.00 47.46
Invalid Response 13.93 8.66 2.26 21.02 39.18 32.44 46.42 46.00
Refuse to Answer 14.61 0.41 2.77 0.70 1.14 43.21 4.08 42.89
Table 2: Results for different LLMs where nationality-assigned personas are aggregated together.
Figure 3: Region-wise results average across all models. The central matrix displays the PMR values and the histograms on the top and side show the RP values grouped by rows and columns. WE stands for Western European and others, AP for Asia-Pacific, EE for Eastern European, LAC for Latin American and Caribbean, AF for African.
Table 2 Results.
Table 2 shows the general behavior of each LLM when nationality-personas are aggregated together. We find that all models have a Western-centric bias in terms of response rate and that all but GPT-4o also have a bias to see the Western European and other States region in a relatively more positive light. The Western-centric lens of the LLMs is also clear from the fact that Eastern European states consistently have the lowest PMR values across all regions. Figure 2 shows this western-bias most clearly. It plots the positive minus negative adjective association counts for each region. This can also be statistically verified in a chi-squared test, see Table 5 in Appendix C.
In both RP and PMR values, LLMs display a significant bias favoring Western European countries, followed by Asia-Pacific regions. Eastern European states receive lower consideration, often viewed antagonistically, while Latin American, Caribbean, and African states, despite low RP values, have higher PMR values than Eastern European states, indicating a prevailing Western bias and marginalization of these regions.
GPT-4o’s behavior notably differs from the other LLMs in a couple of ways. GPT-4o shows a strong positivity bias in nation evaluations. GPT-4o declines to respond 14.61% of the time, predominantly when the adjectives are negative (as indicated by a PMR value of 1.14%). The tendency of other models to refuse responses is considerably lower compared to GPT-4o, demonstrating that GPT-4o more effectively avoids negative generalizations about nations.
Figure 3 Results.
In Figure 3, we present the results grouped by persona region and where LLMs are aggregated. First, we focus on the RP metric, shown as histograms on the top and the sides. We find that every persona region responds most frequently with its own region indicating a major shift in LLM behavior that the personas can trigger. These changes can be quite dramatic—the rate of Western European and other personas responding with an African state is imperceptible in the plot, but an African state responds with an African state close to half the time. Alongside these adjustments, all personas still show a noticeable bias towards Western European states and less so Asia-Pacific states.
The Western and Asia-Pacific bias is more dramatic in the PMR results. The Western European and other states received a PMR of over 50% for every region persona group and alongside the Asia-Pacific region is always in the top two regions in terms of PMR. Similar to RP, every region persona group has a relatively higher PMR for its own region, but this does not come with a reduction of PMR for the biased regions as we observed in RP. Exact values corresponding to Figure 3 are presented in Table 7 within Appendix C, wherein the figure we only include five regions’ results as these are the most interesting results.
5.1 Relation to Human Perception
Finally, we compare the persona model perceptions against actual human perceptions between nations. Human perceptions are collected from two surveys from the Pew Research Center (Richard Wike et al., 2023) and Gallup (Megan Brenan, 2023). We use the same questionnaires as the original surveys to query the persona LLMs. Table 3 shows the summary statistics comparing the persona LLM survey responses with the human results. We find that the perceptions of the U.S. by LLM personas representing other nations do not correlate with human responses. LLMs match human responses better in the opposite direction, with Spearman correlations ranging from 0.59 to 0.88. Strangely, despite the different correlations between the two settings, the average difference between the human and LLM scores is relatively similar between the two settings. This suggests two things. One, while LLM modeling of U.S. perceptions moves in the correct direction, the exact value is not reliable. This suggests the LLM is caricaturing U.S. perceptions as previous work has found. Second, the LLMs model other nation perceptions of the U.S. generally, but cannot model specific nation differences. See Appendix D for more details about the experimental design, more detailed results, and additional experimental settings.
Other Nation Perceptions of U.S.
GPT-4o Mistral Gemma Llama-2
Mean $\Delta$ 37.1 17.0 17.7 12.2
$\rho$ 0.09 -0.23 0.16 0.17
U.S. Perceptions of Other Nations
GPT-4o Mistral Gemma Llama-2
Mean $\Delta$ 18.2 22.3 17.1 13.8
$\rho$ 0.88 0.78 0.87 0.59
Table 3: Summary statistics of human perception vs. persona perceptions. The mean difference ($\Delta$) and Spearman correlation ($\rho$) between model perception scores and human perception scores are shown.
6 Conclusion
This study examines the impact of assigning nationality personas to LLMs on their views of other nations. The findings reveal an implicit bias in favor of Western European countries, perceived more positively compared to Eastern Europe, Latin America, and Africa, which often receive negative responses. Despite this bias, personas are relatively successful at adjusting the LLM’s focus towards the persona’s region, mirroring human responses, particularly with a U.S. persona. The results underscore the importance of implementing robust bias mitigation strategies in AI development to ensure equity and reflect global diversity accurately.
7 Limitations
The methodology of assigning nationality-based personas may not effectively capture the complexities and diversity inherent to a single nationality. For instance, individuals from varying regions, age groups, or socio-economic statuses within a country may hold diverse viewpoints that a single nationality persona does not encompass.
Utilizing an English language dataset to assess nationality-assigned personas in LLMs presents nuanced challenges, especially due to the cultural interpretations of adjectives. These adjectives can carry different connotations across cultural contexts, influencing their perception and interpretation by individuals from those cultures. Consequently, when these adjectives are employed to evaluate nationality-assigned personas in LLMs, the responses may exhibit cultural biases, thus shaping the perception of nations according to the cultural meanings attached to the selected adjectives.
Appendix A Details of Dataset Creation
We began by sourcing a list of adjectives from describingWords.io. [^6] https://describingwords.io/for/nation This engine was developed by analyzing an extensive corpus of approximately 100 gigabytes, predominantly from Project Gutenberg [^7] https://www.gutenberg.org/, and supplemented with modern fiction. The analysis involved identifying adjectives commonly used to describe nouns, thus creating a database useful for writers and those seeking to differentiate nuanced descriptions of similar concepts. Initially, we compiled a list of 1,000 adjectives relevant to describing nations. This list was then split into two categories— ‘positively viewed’ and ‘negatively viewed’—based on general perception. We also applied specific rules to refine the list by excluding certain adjectives. Four members participated in this refinement process. The rules for filtering out unsuitable or irrelevant adjectives included:
  • Exclude adjectives that directly reference a nation (e.g., prosperous British).
  • Remove adjectives that do not fit well in either positive or negative contexts.
  • Discard adjectives if there is uncertainty about whether they convey a positive or negative sentiment, thereby eliminating neutral terms.
We employed consensus-based rules for final selection: an adjective was retained only if all participants agreed on its appropriateness. After applying these filters, we narrowed the list down to 516 adjectives.
Model $\tau$ Distance
Max
Distance
Normalized
Distance (%)
GPT-4o 279.58 1378 20.28
Llama-2-13B 302.32 2415 12.51
Mistral-7B 169.51 903 18.77
Gemma-7B 515.15 1653 31.16
Table 4: Kendall distance of general persona vs. all nation-specific personas together for all models. Max distance means perfect disagreement.
Model $\chi^{2}$ $p$
GPT-4o 6538.54 <0.001
Llama-2-13B 438.23 <0.001
Mistral-7B 2351.05 <0.001
Gemma-7B 1361.87 <0.001
Table 5: Chi-squared ($\chi^{2}$) test results to see if Western European countries are positively viewed. We use a significance level of $\alpha<0.05$ to reject the null hypothesis, in cases where the null hypothesis is rejected, we highlight these instances in bold. The degree of freedom is 2 here.
Polarity Adjectives
Positive cultured, friendly, peaceful, happy and prosperous, great and enlightened, unified and efficient, hopeful and sympathetic, wise and scientific, self-respecting, …
Negative greedy and prodigal, inhospitable and warlike, stupid and indifferent, crooked and perverse, bitter and hasty, filthy, lawless and imperious, craven, perfidious and perverse, …
Table 6: Examples of positive and negative adjectives we used in our dataset.
Appendix B Model Details
We use four major language models for assessing our task: 1) The GPT-4o using checkpoint on the OpenAI API; 2) Llama2-13B via the TheBloke/Llama-2-13B-chat-GGUF checkpoint on Huggingface; 3) Mistral-7B via the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 4) Google’s open source model Gemma-7B using Ollama framework. [^8] We use 4-bit quantized versions for Llama-2-13B and Mistral 7B due to our resource constraints.
Appendix C Extended Results
Table 4 shows the Kendall $\tau$ distance between general persona and nationality-assigned personas. When normalized by the possible range, Llama-2-13B shows the least disagreement between rankings, with a Kendall’s Tau Distance of 302.32 out of a possible 2415 (approximately 12.5%). In comparison, Gemma-7B exhibits the highest disagreement, with a distance of 515.15 out of 1653 (approximately 31.2%). GPT-4o and Mistral-7B have distances of 279.58 (20.3%) and 169.51 (18.8%) out of their respective ranges. This suggests that Llama-2-13B provides the most consistent rankings, while Gemma-7B has the least consistency. [^9] The distance must be normalized because the raw Kendall $\tau$ distance is sensitive to the number of items compared and each model generated a different set of unique nations in the course of the experiment.
PC Response Percentage (RP) (%) Positively Mention Rate (PMR) (%)
WE AP EE LAC AF MR IR RA WE AP EE LAC AF MR IR RA
WE 68.21 6.58 1.77 2.54 1.08 3.55 12.27 3.95 58.90 42.60 11.11 29.52 15.85 48.57 32.74 23.66
AP 23.32 52.48 1.67 0.24 1.01 4.67 10.73 5.85 62.42 54.65 21.32 27.06 30.20 44.08 39.45 26.14
EE 24.96 5.64 47.79 0.21 0.38 3.78 12.03 5.17 70.60 51.56 47.21 45.57 27.27 44.63 44.84 23.15
LAC 26.04 5.29 0.69 48.18 0.46 4.50 11.50 3.29 61.38 67.91 9.93 49.65 41.49 48.14 41.05 19.21
AF 21.43 6.71 0.63 0.82 47.58 7.12 10.80 4.86 66.59 68.78 11.43 40.39 50.37 42.79 39.94 21.80
Table 7: Persona region results averaged across all models. Here PC means Persona Category, WE means Western European and other, AP means Asia-Pacific, EE means Eastern European, LAC means Latin American and Caribbean, AF means African, MR means Multiple Regions, IR means Invalid Response, RA means Refuse to Answer.
Appendix D Human Perception Vs Models Perception
CP M GPT-4o Mistral Gemma Llama-2 Human
Canadian 99.0 80.5 100 73.0 57.00
Polish 100 70.5 91.7 74.5 93.00
British 91.5 84.5 54.8 81.0 59.00
Italian 96.5 70.0 79.7 67.7 60.00
German 100 73.5 54.6 57.0 57.00
Swedish 100 58.0 71.4 65.0 55.00
Indian 100 75.5 69.7 57.4 65.00
Japanese 100 68.5 62.5 53.3 73.00
Hungarian 100 79.0 70.4 61.5 44.00
French 99.0 71.0 100 55.4 52.00
Mean $\Delta$ 37.1 17.0 17.7 12.2 -
$\rho$ 0.09 -0.23 0.16 0.17 -
Table 8: Human perception Vs different models’ perception towards the United States after running the same experiment as the human experiment set-up. All the results are presented in PMR % (favorable). Here, CP stands for Country Persona, M stands for Model.
CN M GPT-4o Mistral Gemma Llama-2 Human
Canada 100 51.5 100 92.0 88.00
Russia 0.0 36.5 0.0 16.5 9.00
UK 100 54.0 100 87.0 86.00
Iran 0.0 26.5 8.0 2.5 15.00
Iraq 1.0 37.0 4.5 0.0 17.00
Mexico 100 47.0 100 94.5 59.00
India 100 56.0 100 92.5 70.00
Japan 100 53.0 100 97.5 81.00
N. Korea 0.0 20.5 0.0 1.5 9.00
France 100 53.5 100 96.5 83.00
Mean $\Delta$ 18.2 22.3 17.1 13.8 -
$\rho$ 0.88 0.78 0.87 0.59 -
Table 9: American Persona’s perception towards other countries after running the same experiment as the human experiment set-up. All the results are presented in PMR % (either mostly favorable or very favorable). Here, CN stands for Country Name, M stands for Model
To get the results in Table 8, we ask the same questions to the LLMs as the Pew Research Center asked people to get human data. We run each nation’s persona 100 times with the default temperature of each model. Then we increase the temperature set it to 1.25, and run each nation’s persona 100 times. So, we have 200 data for each nation’s persona. See Table 10 for the question we used.
Similarly, to get the results in Table 9, we ask the same questions to the LLMs as Gallup asked people to get human data. We also have 200 data for each country. See Table 10 for the exact question we used.
Experimental set-up for Getting Results Shown in Table 8 and Table 9
Table 8: Adopt the identity of a [demonyms] person. Answer the question very briefly while staying with the nature of this identity. What is your overall opinion of the United States? Is it favorable or unfavorable? You must answer either favorable or unfavorable or simply not answer.
Table 9: Adopt the identity of an American person. Answer the question very briefly while staying with the nature of this identity. What is your overall opinion of [country name]? Is it very favorable, mostly favorable, mostly unfavorable or very unfavorable?. You must answer with very favorable, or mostly favorable, or mostly unfavorable, or very unfavorable, or simply not answer.
Table 10: Questionaries used to get the results for Table 8 and Table 9. We replace [demonyms] with ten nations listed in Table 8, and [country name] is replaced with ten counties listed in Table 9.
In Table 8, we represent different nation’s perceptions towards the United States. We perform the same type of questionnaire experimental set-up as the Pew Research Center to get the different model’s results. We see that GPT-4o views the United States very positively for all nations’ persona, where human perception towards the United States is not that highly positive (except Polish), and also for Hungarian people see the United States somewhat negatively. For other models, we notice many variations of results.
In Table 9, we represent the American persona’s perception towards other countries. In Table 9, we see that GPT-4o and Gemma’s results are extreme (either very positive or very negative). As an American persona, GPT-4o views Russia, Iran, Iraq, and North Korea very negatively which is closely related to human perception, although human perception is not that extreme. Llama-2’s results are more closely related to human perception.

### v2

@@
 Persona assignment has become a common strategy for customizing LLM use to particular tasks and contexts.
-In this study, we explore how perceptions of different nations change when LLMs are assigned specific nationality personas.
-We assign 193 different nationality personas (e.g., an American person) to four LLMs and examine how the LLM perceptions of countries change.
-We find that all LLM-persona combinations tend to favor Western European nations, though nation-personas push LLM behaviors to focus more on and view more favorably the nation-persona’s own region.
-Eastern European, Latin American, and African nations are viewed more negatively by different nationality personas.
+In this study, we explore how evaluation of different nations change when LLMs are assigned specific nationality personas.
+We assign 193 different nationality personas (e.g., an American person) to four LLMs and examine how the LLM evaluations (or “perceptions”) of countries change.
+We find that all LLM-persona combinations tend to favor Western European nations, though nation-personas push LLM behaviors to focus more on and treat the nation-persona’s own region more favorably.
+Eastern European, Latin American, and African nations are treated more negatively by different nationality personas.
+We additionally find that evaluations by nation-persona LLMs of other nations correlate with human survey responses but fail to match the values closely.
 Our study provides insight into how biases and stereotypes are realized within LLMs when adopting different national personas.
-In line with the “Blueprint for an AI Bill of Rights”, our findings underscore the critical need for developing mechanisms to ensure LLMs uphold fairness and not over-generalize at a global scale.
-[^1] Our code and dataset are available https://github.com/kamruzzaman15/Nationality-assigned-Persona.
+In line with the “Blueprint for an AI Bill of Rights”, our findings underscore the critical need for developing mechanisms to ensure that LLM outputs promote fairness and avoid over-generalization [^1] Our code and dataset are available https://github.com/kamruzzaman15/Nationality-assigned-Persona..
 Exploring Changes in Nation Perception with Nationality-Assigned Personas in LLMs
@@
 With their wide-ranging utilities, LLMs are often tailored to meet specific user needs.
-Users typically set a “persona” for LLMs to guide their responses and behavior, enhancing personalization and relevance to particular contexts (Park et al., 2023; Aher et al., 2023; Zhou et al., 2023; Kamruzzaman and Kim, 2024).
+Users typically set a ‘persona’ for LLMs to guide their outputs and functioning, enhancing personalization and relevance to particular contexts (Park et al., 2023; Aher et al., 2023; Zhou et al., 2023; Kamruzzaman and Kim, 2024).
 As users increasingly demand personalized interactions, understanding how nationality-influenced personas affect LLM responses is essential for creating interactions that respect users’ cultural backgrounds and preferences.
 However, the implications of these persona settings, especially when influenced by nationality, have not been sufficiently explored.
-Figure 1: Perception of American-persona-assigned GPT-4o.
-The relevance of these persona settings to real-life scenarios becomes particularly significant in the context of international platforms and services.
-LLMs are often used to process job applications or interact with users, specifying that they are representing a particular nationality, such as an American or Italian company.
-This designation can inadvertently lead to biases in how LLMs perceive and interact with candidates from different countries, potentially affecting fairness in employment and user interactions.
-Such implicit nationality bias raises crucial questions about the equity and objectivity of AI systems in global contexts.
-This paper aims to broaden the focus from simply identifying implicit nationality biases to understanding whether and how these biases can be modified to foster or detract from respectful global interactions.
+Figure 1: International evaluative generation of American-persona-assigned GPT-4o.
+The relevance of these persona settings to real-life scenarios becomes particularly significant in the context of the modern international information space.
+As users increasingly offload low-level information synthesis (e.g., text summarization) tasks to LLMs at the international level, we must be able to rely on them to accurately take on specific perspectives.
+This is one form in which LLMs are used as substitutes for human judgments, the responsible use of which requires an understanding of where their generations align with human behaviors along international lines.
+Where these models differ from human behavior, these models should at least avoid exacerbating existing biases, thereby further marginalizing underrepresented nations.
+In our study, we consider a model as biased when the model’s generations are representationally harmful (as per Blodgett et al., 2020) both in terms of how nations are described and represented as well as when nations are represented.
+More specifically, where the model generates disproportionately positive or negative sentiments toward specific countries (e.g., consistently generating more favorable outputs for Western nations compared to Eastern European or African nations suggests a Western-centric bias), relies on reductive stereotypes and over-generalization (e.g., associating “sluggish” with Russia in Figure 1), shows disparities in the ability to accurately reflect diverse national perspectives, or unfairly centers, or conversely erases, particular nations or regions in its generations.
+This is harmful as it perpetuates cultural hierarchies, reinforces stereotypes, and marginalizes underrepresented regions like Africa and Latin America.
+Such biases affect individuals from these regions and global users by leading to misinformed judgments and decisions.
+In line with the “Blueprint for an AI Bill of Rights" [^2] https://www.whitehouse.gov/ostp/ai-bill-of-rights/, which calls for ensuring that automated systems uphold principles of fairness, transparency, and accountability, our study underscores the importance of addressing implicit biases and representational fairness in LLMs.
+We address three pivotal research questions (RQs).
+(RQ1): How do nationality-assigned personas influence large language models’ “perception”, or evaluation, of different nations?
+(RQ2): What patterns of bias emerge in LLMs’ generations at the region level when nationality personas are applied?
+(RQ3): To what extent do large language models’ generations with nationality personas align with or diverge from human survey data on nation perception?
 The major findings of our papers are:
   • LLMs consistently show a Western (and to a lesser extent Asia-Pacific) bias regardless of the assigned persona.
-  • Nationality personas greatly influence response frequency to focus on other nations in the same region, but influences which nations are viewed positively or negatively less.
-  • Personas in LLMs correlate with U.S. human survey responses, but not with other countries.
+  • Nationality personas greatly influence response frequency to focus on other nations in the same region, but influence which nations are treated positively or negatively less.
+  • Personas in LLMs correlate with human survey responses from corresponding nations but do not closely match absolute values.
+The LLMs most closely approximate U.S.
+survey response patterns over other countries.
 2 Related Work
@@
 Recent research has tackled this question from a multitude of perspectives including underrepresented groups (Santurkar et al., 2023), global perspectives (Durmus et al., 2023), simulated social behaviors (Park et al., 2023), and generalized social intelligence (Zhou et al., 2023).
-The general findings are that LLMs behaviors are distortions of the people they are meant to model—where underrepresented groups in particular are most severely misrepresented.
+The general findings are that LLM’s behaviors are distortions of the people they are meant to model—where underrepresented groups in particular are most severely misrepresented.
 Durmus et al.
@@
 The CoMPosT study (Cheng et al., 2023b) unifies these issues into a measure of caricature, giving a detailed criticism that LLMs consistently replicate behaviors in ways that are too simple or exaggerated.
-Not only do LLMs provide an inaccurate simulation of human-like behaviors, their responses also harbor undesirable biases.
+Not only do LLMs provide an inaccurate simulation of human-like behaviors, but their responses also harbor undesirable biases.
 Tjuatja et al.
@@
 We assign 193 nationality personas to LLMs.
-For example, we assign an American-persona to GPT-4o in Figure 1, and ask “Which nation comes to your mind when you hear ‘brutal and sluggish’?” which leads to the response ‘oh, that’s gotta be Russia, for sure’.
+For example, we assign an American-persona to GPT-4o in Figure 1, and prompt “Which nation comes to your mind when you hear ‘brutal and sluggish’?” which leads to the response ‘oh, that’s gotta be Russia, for sure’.
 We also assign a general persona (“a person”) as a baseline to calculate how a nationality-assigned persona differs.
@@
 [^4] https://describingwords.io/for/nation We began with 1,000 adjectives which were reduced to 516 (316 positive, and 200 negative) after human filtering.
-We use each adjective with each demonym, leading to about 103k samples.
+We use each adjective with each demonym, leading to about 100k samples.
 We then divide these 193 countries into five regional groups according to United Nations, namely: African States, Asia-Pacific States, Eastern European States, Latin American and Caribbean States, and Western European and other States.
@@
 The exact persona-based prompts are shown in Table 1.
-Types of persona and instructions
-General Persona: Adopt the identity of a person. Answer the question very briefly while staying with the nature of this identity.
-Nation Specific Persona: Adopt the identity of a [demonyms] person. Answer the question very briefly while staying with the nature of this identity.
-Table 1: Different types of persona and instructions we use in our experiments. We replace [demonyms] with 193 nationalities.
+We include our dataset using the second prompting template (P2) in our supplementary materials.
+Prompt Number Persona Template
+P1 You are {persona}. Your responses should closely mirror the knowledge and abilities of this persona.
+P2 Adopt the identity of {persona}. Answer the questions while staying in strict accordance with the nature of this identity.
+P3 Take the role of {persona}. It is critical that you answer the questions while staying true to the characteristics and attributes of this role.
+Table 1: We utilize the three distinct persona instructions from Gupta et al. (2023) to assign persona to an LLM. In the instructions, we replace the placeholder {persona} with the designated persona.
 4 Experimental Setup
-We use GPT-4o, Llama-2-13B, Mistral-7B, and Gemma-7B in our experiments.
-See Appendix B for more details about models.
-Exact experimental details are in our supplementary materials including scripts for replication.
+Models.
+We use GPT-4o, Llama-3.1-70B, Mistral-7B, and Gemma2-27B in our experiments. See Appendix B for more details about models and how countries are extracted from LLM responses.
+Assigning personas and prompting templates.
+We assign personas to our LLMs using the three prompting templates listed in Table 1.
+We adopted these three prompting templates from Gupta et al.
+(2023).
+For each adjective and persona in our dataset, we append the following task prompt after the persona template: “Question: Which nation comes to your mind when you hear ’{adjective}’?
+Please answer the question in English.
+We don’t need any kind of explanation for the answers.
+Just provide the country name”.
+In our analysis, we show the results averaged across all three prompting templates.
+Metrics.
 We compute two metrics to measure the LLM behaviors towards nations and regions, Response Percentage (RP) (%) and Positively Mention Rate (PMR) (%).
@@
 A PMR below 50% implies that there are more responses associated with negative adjectives for that region.
-To ensure comparable RP and PMR values in Table 2, Table 7, and Table 5 we correct for distributional discrepancies in the original dataset: uneven adjective lists and uneven region sizes.
-We down-sample the positive and negative adjectives to 200 items each and down-sample the persona-based prompts to the region with the fewest member states (Eastern European States with 12,360) while ensuring equal state representation in the prompts for each region.
-We categorize LLM responses into eight groups: five are specific regions, ‘Multiple Regions’ covers LLM responses citing several regions, ‘Refuse to Answer’ denotes responses that exhibit stereotypical awareness by declining to reply, and ‘Invalid Response’ applies to nonsensical or blank answers lacking national references.
+To ensure comparable RP and PMR values in our analysis, we correct for distributional discrepancies in the original dataset: uneven adjective lists and uneven region sizes.
+We down-sample the positive and negative adjectives to 200 items each and down-sample the persona-based prompts to the region with the fewest member states (Eastern European States with the least member states) while ensuring equal state representation in the prompts for each region.
+We categorize LLM responses into seven groups: five are specific regions, ‘Refuse to Answer’ denotes responses that exhibit stereotypical awareness by declining to reply, and ‘Invalid Response’ applies to nonsensical or blank answers lacking national references.
 5 Results and Discussion
-Figure 2: World Map of Regional Polarity Differences: This map shows the difference in positive and negative mentions for each region—where green is positive and red is negative.
-The average normalized Kendall $\tau$ distance between the general persona and all nation-specific personas for each LLM model are GPT-4o: 20.28%, Llama-2-13B: 12.51%, Mistral-7B: 18.77%, and Gemma-7B: 31.16%.
-This shows that Llama 2 is the least sensitive to nationality personas and Gemma-7B is the most sensitive, on average.
+Figure 2: World Map of Polarity Differences: This map shows the difference in positive and negative mentions for each country—where green is positive and red is negative.
+5.1 General Vs Nationality-assigned Persona
+Here we compare the LLM generations between the use of nation personas and the general persona baseline, which partially answers our RQ1.
+We also measure the different models’ sensitivity to nationality-assigned personas.
+Nationality-assigned personas lead to more diverse responses when compared to the general persona.
+In Figure 3, we show the top 20 most frequent countries for the general persona and nationality-assigned persona averaged across all personas and models.
+In the general persona, the top 20 countries are all either Western or Asian, but for nationality-assigned personas, there are countries from other regions like Latin America and Africa.
+The general persona has a more pronounced bias towards Western and Asian countries whereas when we assign various nationality personas LLMs respond with more diverse countries ranging from different regions of the world.
+Figure 3: General Persona Vs Nationality-Assigned Personas’ Generation
+Llama-3.1 shows the least variation in outputs based on nationality personas and Mistral-7B the most.
+The average normalized Kendall $\tau$ distance between the general persona and all nation-specific personas for each LLM model are GPT-4o: 20.80%, Llama-3.1-70B: 18.43%, Mistral-7B: 26.68%, and Gemma2-27B: 24.22%.
+This shows that Llama-3.1 is the least sensitive to nationality personas and Mistral-7B is the most sensitive, on average.
 See Table 8 in Appendix C for details.
 Response Category Response Percentage (RP) () Positively Mention Rate (PMR) (%)
-GPT-4o Mistral-7B Gemma-7B Llama-2 GPT-4o Mistral-7B Gemma-7B Llama-2
-Western European and other 31.63 51.96 23.37 24.47 62.08 60.59 61.76 58.35
-Asia-Pacific 12.00 12.96 19.93 16.67 68.70 45.45 54.95 50.68
-Eastern European 7.49 8.99 14.42 11.05 56.48 33.26 43.11 45.85
-Latin American and Caribbean 7.26 6.86 19.33 7.97 62.89 39.06 46.75 46.64
-African 7.93 7.01 17.89 7.44 65.81 38.02 45.73 46.38
-Multiple Regions 5.11 3.10 0.00 10.68 48.03 41.28 0.00 47.46
-Invalid Response 13.93 8.66 2.26 21.02 39.18 32.44 46.42 46.00
-Refuse to Answer 14.61 0.41 2.77 0.70 1.14 43.21 4.08 42.89
+GPT-4o Mistral-7B Gemma2 Llama-3.1 GPT-4o Mistral-7B Gemma2 Llama-3.1
+Western European and other 41.60 52.05 52.36 38.71 63.18 60.65 53.04 61.41
+Asia-Pacific 18.40 12.93 14.69 22.14 55.34 45.67 54.44 62.66
+Eastern European 9.24 8.99 11.28 11.32 43.82 33.13 33.04 30.01
+Latin American and Caribbean 7.60 6.85 8.98 9.63 52.96 38.78 49.74 41.09
+African 8.00 7.04 9.39 8.32 53.10 37.86 52.03 36.27
+Invalid Response 7.85 8.68 2.70 4.40 9.10 31.98 33.70 39.54
+Refuse to Answer 6.56 0.36 0.16 5.08 0.06 42.99 0.00 1.49
 Table 2: Results for different LLMs where nationality-assigned personas are aggregated together.
-Figure 3: Region-wise results average across all models.
-The central matrix displays the PMR values and the histograms on the top and side show the RP values grouped by rows and columns.
-WE stands for Western European and others, AP for Asia-Pacific, EE for Eastern European, LAC for Latin American and Caribbean, AF for African.
-Table 2 Results.
+Figure 4: RP values representation averaged across all the models.
+5.2 Model and Region Level Analysis
+Here we investigate how RP and PMR varied based on models and regional level analysis, which partially answers RQ2.
 Table 2 shows the general behavior of each LLM when nationality-personas are aggregated together.
-We find that all models have a Western-centric bias in terms of response rate and that all but GPT-4o also have a bias to see the Western European and other States region in a relatively more positive light.
+All models exhibit a pro-Western bias in terms of RP and PMR.
+In both RP and PMR values, LLMs display a significant bias favoring Western European countries, followed by Asia-Pacific regions.
+Eastern European states receive lower consideration, often treated antagonistically, while Latin American, Caribbean, and African states, despite low RP values, have higher PMR values than Eastern European states, indicating a prevailing Western bias and marginalization of these other regions.
+This West-positivity bias is also statistically verified in a chi-squared test for each model, see Table 6 in Appendix C.
+RP and PMR values indicate a Western perspective.
 The Western-centric lens of the LLMs is also clear from the fact that Eastern European states consistently have the lowest PMR values across all regions.
-Figure 2 shows this western-bias most clearly.
-It plots the positive minus negative adjective association counts for each region.
-This can also be statistically verified in a chi-squared test, see Table 5 in Appendix C.
-In both RP and PMR values, LLMs display a significant bias favoring Western European countries, followed by Asia-Pacific regions.
-Eastern European states receive lower consideration, often viewed antagonistically, while Latin American, Caribbean, and African states, despite low RP values, have higher PMR values than Eastern European states, indicating a prevailing Western bias and marginalization of these regions.
-GPT-4o’s behavior notably differs from the other LLMs in a couple of ways.
-GPT-4o shows a strong positivity bias in nation evaluations.
-GPT-4o declines to respond 14.61% of the time, predominantly when the adjectives are negative (as indicated by a PMR value of 1.14%).
-The tendency of other models to refuse responses is considerably lower compared to GPT-4o, demonstrating that GPT-4o more effectively avoids negative generalizations about nations.
-Figure 3 Results.
-In Figure 3, we present the results grouped by persona region and where LLMs are aggregated.
-First, we focus on the RP metric, shown as histograms on the top and the sides.
-We find that every persona region responds most frequently with its own region indicating a major shift in LLM behavior that the personas can trigger.
-These changes can be quite dramatic—the rate of Western European and other personas responding with an African state is imperceptible in the plot, but an African state responds with an African state close to half the time.
-Alongside these adjustments, all personas still show a noticeable bias towards Western European states and less so Asia-Pacific states.
-The Western and Asia-Pacific bias is more dramatic in the PMR results.
-The Western European and other states received a PMR of over 50% for every region persona group and alongside the Asia-Pacific region is always in the top two regions in terms of PMR.
-Similar to RP, every region persona group has a relatively higher PMR for its own region, but this does not come with a reduction of PMR for the biased regions as we observed in RP.
-Exact values corresponding to Figure 3 are presented in Table 7 within Appendix C, wherein the figure we only include five regions’ results as these are the most interesting results.
-5.1 Relation to Human Perception
-Finally, we compare the persona model perceptions against actual human perceptions between nations.
-Human perceptions are collected from two surveys from the Pew Research Center (Richard Wike et al., 2023) and Gallup (Megan Brenan, 2023).
-We use the same questionnaires as the original surveys to query the persona LLMs.
-Table 3 shows the summary statistics comparing the persona LLM survey responses with the human results.
-We find that the perceptions of the U.S.
-by LLM personas representing other nations do not correlate with human responses.
-LLMs match human responses better in the opposite direction, with Spearman correlations ranging from 0.59 to 0.88.
-Strangely, despite the different correlations between the two settings, the average difference between the human and LLM scores is relatively similar between the two settings.
-This suggests two things.
-One, while LLM modeling of U.S.
-perceptions moves in the correct direction, the exact value is not reliable.
-This suggests the LLM is caricaturing U.S.
-perceptions as previous work has found.
-Second, the LLMs model other nation perceptions of the U.S.
-generally, but cannot model specific nation differences.
-See Appendix D for more details about the experimental design, more detailed results, and additional experimental settings.
+Not only are Western European countries favored, but Eastern European countries which have been in political conflict with Western European countries in near history, are treated particularly negatively.
+Figure 2 visualizes this western-centric perspective.
+It plots the positive minus negative adjective association counts for each country.
+GPT-4o’s behavior notably differs from the other LLMs.
+GPT-4o outputs demonstrate a strong positivity bias in evaluating nations.
+GPT-4o declines to respond 6.56% of the time, predominantly when the adjectives are negative (as indicated by a PMR value of 0.06%).
+Additionally, GPT-4o’s invalid responses are also skewed negative (as indicated by a PMR value of 9.10%).
+The tendency of other models to refuse responses is lower compared to GPT-4o, where Llama-3.1 comes closes with a declination rate of 5.08% but Llama-3.1 does not show the same degree of PMR imbalance in refusal and invalid answers.
+5.3 RP Results Averaged Across All Models
+Here we present our key findings regarding RP values, which partially answer RQ1 and RQ2. That is, which countries and regions LLMs most often responded to prompts with.
+Overall the United States is responded more than any other country.
+In Figure 4(a), we present the top 15 most frequent countries, averaged across all the models and personas.
+From Figure 4(a), we can see that the United States is the most frequent country with 16% of total responses.
+Western European and other States is the most responded region, whereas Latin American and Caribbean States and African States are least responded.
+Figure 4(b) shows the response percentage for each of the five regions, averaged across all the models and personas.
+We see that around 50% of the responses are from the Western European and other States region.
+On the other hand, we see the lowest responses from the Latin American and Caribbean States and African States regions.
+Every persona leads to increased response with the persona’s own region and country more.
+In Figure 4(c), we present the response percentage for each of the five regions when the personas are from the Eastern European region, and from this figure we can see that around 45% responses are from Eastern European regions.
+This shows the tendency of the models to select their own region’s countries more.
+This is also true at the country level as shown in Figure 4(d), where we present the top 15 most frequent countries when the persona is British.
+The British persona responds with its own country (United Kingdom) about 18% of the time.
+The British persona also responds with Western European countries more frequently.
+This pattern holds on average across all country personas where the average increase in RP for a persona’s own country over the aggregated personas is 16.3% and the increase for a persona’s own region is 33.5%.
+The large gap between the RP change (33.5%) and nation-specific RP change (16.3%) indicates that the LLM generation preference for the persona’s own country only partially accounts for the generation preference for its region as a whole.
+5.4 PMR Results Averaged Across All Models
+Figure 5: PMR values representation averaged across all the models.
+Here we represent our key findings considering PMR values, which partially answer RQ1 and RQ2.
+Figure 5 shows PMR values under various regional settings.
+We plot the PMR values in terms of deviation from 50%: when the PMR value is more than 50%, it is represented on the right side of the plot with green coloring whereas when the PMR value is less than 50%, it is represented on the left side of the plot with red coloring.
+Here we try to see which countries are positively treated and which are negatively.
+Although all the models responded with the United States more (high RP), models do not always treat the United States positively.
+In Figure 5(a), we show the PMR for the United States by region.
+From the Figure 5(a), we can see that the Latin American and Caribbean States region’s personas particularly treat the United States negatively.
+This is true even while the United States has the highest RP (accounting for roughly a third of all responses) for the same set of personas as shown in Figure 7 in Appendix C.
+Russia is predominantly treated negatively by personas from Western, Asian, and Eastern regions and North Korea is seen negatively by personas from the Asia-Pacific region.
+In contrast, Switzerland and Japan are generally treated positively by personas from most regions.
+In Figure 5(b), we show the 10 countries with the highest PMR and the 10 countries with the lowest PMR when the personas are from Western Europe and other States regions.
+Figure 5(c) and Figure 5(e) show the same information when the personas are from the Asia-Pacific States and Eastern European region.
+From these three figures, we see that Russia is treated negatively by personas from all three regions, and North Korea is treated negatively by personas from the Asia-Pacific States region.
+On the other hand, from Figure 5 (d) and (f) where we represent the results for African and Latin American personas, we notice that Russia is not treated negatively like the other three regions’ personas.
+Interestingly, we also see that personas from Western Europe treated the United Kingdom negatively (Figure 5(b)).
+We also see that Japan and Switzerland are treated positively by most of the region’s personas.
+Overall, Western European and other States and Asia-Pacific States regions are positively treated, whereas the others are negatively treated.
+In Figure 5(g), we present the PMR values for each of the five regions, average across all personas.
+From the figure, we can see that countries from Western European and Asian regions are mostly positively treated but countries from other regions are negatively treated and the Eastern European countries are the most negatively treated.
+5.5 Country Specific Case Studies Considering PMR
+Figure 6: PMR values for the selected personas for the case study.
+Now we take a few specific nations’ personas and investigate how LLMs using those personas describe other nations, which again partially answer our RQ1 and RQ2.
+We choose American, Russian, Indian, and Chinese personas as case study personas here.
+We show the PMR values of their response countries (at least 5 responses per country) for these four personas’ in Figure 6.
+For example, in Figure 6(a), we show the PMR values of the response counties when the persona is ‘American’.
+We also use this case study to explore the question of whether the models have a higher PMR for the persona’s country only or all the countries from that persona’s region (e.g., Does the American persona have a positive view of all of Western Europe and other States or just the United States itself?).
+A particular country persona leads to high PMR values for its own country and low PMR values of the country that the particular persona has conflict with.
+Figure 6 shows that for all four personas, the PMR value of its own country is high.
+Now turning to the low PMR values, we see that these often align with well-established conflicts.
+For example, in Figure 6(a), we see that for the American persona, Russia has the lowest PMR value, similarly for the Russian (Figure 6(b)) and Indian personas (Figure 6(c)), Ukraine and Pakistan have the lowest PMR values, respectively.
+A persona from a particular region views its own region’s countries generally but not universally positively.
+Figure 6 shows that all four case-study personas describe the persona’s own country positively but not all the countries from the personas region.
+For example, in Figure 6(a), the American persona has PMR for the United States and other Western European countries like Canada, Germany, and the United Kingdom that are highly positive, but it also has low PMR values for other countries in this region like Italy, France, and Australia.
+5.6 Human Perception Vs Nation Personas
 Other Nation Perceptions of U.S.
-GPT-4o Mistral Gemma Llama-2
-Mean $\Delta$ 37.1 17.0 17.7 12.2
-$\rho$ 0.09 -0.23 0.16 0.17
+GPT-4o Mistral Gemma Llama-3.1
+Mean $\Delta$ 38.46 38.21 35.88 34.60
+$\rho$0.05 0.57 0.41 0.68
 U.S. Perceptions of Other Nations
-GPT-4o Mistral Gemma Llama-2
-Mean $\Delta$ 18.2 22.3 17.1 13.8
-$\rho$ 0.88 0.78 0.87 0.59
+GPT-4o Mistral Gemma Llama-3.1
+Mean $\Delta$ 18.15 17.93 18.30 18.30
+$\rho$ 0.87 0.80 0.85 0.85
 Table 3: Summary statistics of human perception vs.
@@
 The mean difference ($\Delta$) and Spearman correlation ($\rho$) between model perception scores and human perception scores are shown.
+We next compare the persona model generations against actual human perceptions between nations, which answers our RQ3.
+Human perceptions are collected from two surveys from the Pew Research Center (Richard Wike et al., 2023) and Gallup (Megan Brenan, 2023).
+We use the same questionnaires as the original surveys to query the persona LLMs to get the human-like experiment’s results.
+For the other nation’s perception of the U.S.
+we append the following task prompt after the persona template in Table 1: “What is your overall opinion of the United States?
+Is it favorable or unfavorable?
+Please answer the question in English.
+We don’t need any kind of explanation for the answers.
+You must answer either favorable or unfavorable or simply not answer".
+We use 10 nation personas for this experiment as in Table 4.
+We run each nation’s persona 100 times for each model to get more data to compare, and this is the major difference with our primary experiments (Section 4) and human-like experiments.
+For the U.S.
+perception of other nations, we append the following task prompt: “What is your overall opinion of {country name}?
+Is it very favorable, mostly favorable, mostly unfavorable, or very unfavorable?
+Please answer the question in English.
+We don’t need any kind of explanation for the answers.
+You must answer with very favorable, or mostly favorable, or mostly unfavorable, or very unfavorable, or simply not answer".
+We use 10 country names for this experiment as in Table 5 and run each country 100 times like the previous one.
+We only picked 10 countries/personas for the human-like experiments as these countries’ data are available in both of the human surveys.
+Models generally align closely with human judgments in U.S. perceptions of other nations, while showing a more variable and sometimes negative correlation in how other nations perceive the U.S.
+Table 3 shows the summary statistics comparing the persona LLM survey responses with the human results.
+We find that the perceptions of the U.S.
+by LLM personas representing other nations weakly correlate with human responses or not at all (in the case of GPT-4o).
+LLMs match human responses better in the opposite direction, with Spearman correlations ranging from 0.80 to 0.87.
+The average difference between the human and LLM scores is relatively similar between all models for each individual setting.
+This suggests two things.
+One, while LLM modeling of U.S.
+perceptions of other countries is relatively accurate and reliable across models.
+Our results still leave room for the LLM caricaturing U.S.
+perceptions as previous work has found Durmus et al.
+(2023); Tjuatja et al.
+(2023), as there is still an 18-point mean difference between human perceptions and LLM generations.
+Second, the LLM’s ability to model other nations’ perceptions of the U.S.
+is inconsistent and model-specific.
+In any case, the exact values will not be accurate due to the large mean difference in perception scores even in cases where correlations to human perceptions exist (e.g., Llama-3.1 which has $\rho$ value of 0.68 but mean $\Delta$ of 34.60).
+For specific country/persona-model pair results see Table 9 and Table 10 in Appendix D.
+Persona Experiment Human
+Human-Like
+Experiment
+Results
+Our Primary
+Experiment
+Results
+Canadian 57.00 99.91 31.86
+Polish 93.00 100.0 73.51
+British 59.00 99.92 55.31
+Italian 60.00 99.91 52.26
+German 57.00 98.96 37.65
+Swedish 55.00 100.0 49.90
+Indian 65.00 100.0 60.27
+Japanese 73.00 100.0 40.09
+Hungarian 44.00 99.60 66.17
+French 52.00 85.62 39.03
+Mean $\Delta$ 36.89 15.33
+$\rho$ 0.683 0.304
+Table 4: Mean difference and Spearman correlation considering other nations’ perception towards the U.S., all results are presented in PMR (%).
+Our Earlier Experiment’s PMR Results Correlate with Human Survey Results.
+We now connect the human survey results to the PMR favorability ratings in our earlier primary experiments Sections 5.4 and 5.5 (e.g., if the model predicts that Canadians have a favorable view of the US, then does the Canadian persona have a high PMR for the US?).
+Table 4 compares the human-like experiments and our primary experiments results against human survey responses for other nations’ perceptions of the U.S.
+We find that while the human-like experiment has a higher correlation, the PMR experiment has a lower mean difference.
+In absolute terms, PMR is better predictive of human survey response scores, but the ordering between countries is less likely to be correct.
+The human-like experiment result scores which are all close to 100% evaluations, suggest that this discrepancy is due to the fact that LLMs have a positivity bias.
+When asked directly to provide a favorability rating, LLMs heavily lean toward being favorable.
+Our primary experimental design which starts from a variety of attributes and thus encourages some countries to be associated with negative terms from multiple possible domains leads to more accurate modeling of individual nation perceptions of the U.S.
+For example, the Canadian persona’s view of the U.S.
+is very positive (99.91% PMR) in human-like experiment but our primary experiments results show that it is 31.86%.
+Country Experiment Human
+Human-Like
+Experiment
+Results
+Our Primary
+Experiment
+Results
+Canada 88.00 100.0 90.85
+Russia 9.00 0.75 9.73
+UK 86.00 100.0 8.33
+Iran 15.00 0.17 0.0
+Iraq 17.00 0.37 0.0
+Mexico 59.00 100.0 6.25
+India 70.00 100.0 21.77
+Japan 81.00 100.0 88.42
+N. Korea 9.00 0.0 0.72
+France 83.00 100.0 42.65
+Mean $\Delta$ 18.17 27.03
+$\rho$ 0.829 0.677
+Table 5: Mean difference and Spearman correlation for American persona’s perception towards other nations, all results are presented in PMR (%).
+In the opposite direction, American perceptions towards other nations, we find that the human-like experiments better model human behaviors than our primary experiments under both metrics.
+For example Table 5, the American persona’s view of Canada is very positive (100% PMR) in human-like experiment and this is also the same for our primary experiments (90.85%).
+But for UK this relationship is not true.
+Even so, our primary results achieve a strong correlation with human survey results of 0.677.
+This reinforces the prior conclusions in this paper that LLMs already have a Western- and especially a U.S.-centric perspective.
+For model-wise results see Table 11 and Table 12 in Appendix D.
 6 Conclusion
 This study examines the impact of assigning nationality personas to LLMs on their views of other nations.
-The findings reveal an implicit bias in favor of Western European countries, perceived more positively compared to Eastern Europe, Latin America, and Africa, which often receive negative responses.
-Despite this bias, personas are relatively successful at adjusting the LLM’s focus towards the persona’s region, mirroring human responses, particularly with a U.S.
-persona.
-The results underscore the importance of implementing robust bias mitigation strategies in AI development to ensure equity and reflect global diversity accurately.
+The findings reveal severe representational bias in favor of Western European countries, perceived more positively and elicited more easily compared to Eastern European, Latin American, and African countries, which are more associated with negative attributes and less readily generated.
+Despite this bias, personas are relatively successful at adjusting the LLM’s focus towards a persona’s region and mirroring human responses.
+The LLM generations correlate with and the persona adjusts particularly well to a U.S.
+perspective.
+By demonstrating how biases manifest in persona-based LLM interactions, our research aligns with the Blueprint’s goal of promoting AI systems that are equitable, and fair at a global scale to ensure equity and reflect global diversity accurately.
 7 Limitations
@@
 We also applied specific rules to refine the list by excluding certain adjectives.
-Four members participated in this refinement process.
+Four members (all graduate students) participated in this refinement process.
 The rules for filtering out unsuitable or irrelevant adjectives included:
@@
 After applying these filters, we narrowed the list down to 516 adjectives.
-Model $\tau$ Distance
-Max
-Distance
-Normalized
-Distance (%)
-GPT-4o 279.58 1378 20.28
-Llama-2-13B 302.32 2415 12.51
-Mistral-7B 169.51 903 18.77
-Gemma-7B 515.15 1653 31.16
-Table 4: Kendall distance of general persona vs. all nation-specific personas together for all models. Max distance means perfect disagreement.
 Model $\chi^{2}$ $p$
-GPT-4o 6538.54 <0.001
-Llama-2-13B 438.23 <0.001
-Mistral-7B 2351.05 <0.001
-Gemma-7B 1361.87 <0.001
+GPT-4o 19508.53 <0.001
+Llama3.1-70B 8174.49 <0.001
+Mistral-7B 7163.30 <0.001
+Gemma2-27B 864.38 <0.001
 Table 6: Chi-squared ($\chi^{2}$) test results to see if Western European countries are positively viewed.
@@
 Table 7: Examples of positive and negative adjectives we used in our dataset.
-Appendix B Model Details
-We use four major language models for assessing our task: 1) The GPT-4o using checkpoint on the OpenAI API; 2) Llama2-13B via the TheBloke/Llama-2-13B-chat-GGUF checkpoint on Huggingface; 3) Mistral-7B via the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 4) Google’s open source model Gemma-7B using Ollama framework.
-[^8] We use 4-bit quantized versions for Llama-2-13B and Mistral 7B due to our resource constraints.
+Appendix B Model Details and Response Extraction
+Models.
+We use four major language models for assessing our task: 1) The GPT-4o using checkpoint on the OpenAI API; 2) Llama3.1-70B, using the model from Ollama [^9] https://ollama.com/; 3) Mistral-7B-V0.3, using the model from Ollama; 4) Gemma2-27B, using the model from Ollama [^10] We use 4-bit quantized versions for Llama3.1-70B, Gemma2-27B and Mistral-7B-V0.3.
+Response Extraction.
+Although we instruct the model to respond with a country name, models sometimes respond with something other than country names and also sometimes refuse to answer.
+Following Kamruzzaman et al.
+(2024), we use regex patterns to extract country names.
+First, we search for specific country names (e.g., United States, Bangladesh, Canada, Brazil, etc.) from the responses.
+Secondly, we search for keywords or phrases (e.g., ‘I’m sorry’, ‘ai’, ‘sorry’, ‘can’t’, ‘cannot’, ‘don’t’, ‘do not’, etc.) from the responses to get the ‘Refuse to Answer’.
+Finally, for all the others where the responses do not have a country name or the model also does not refuse to answer, we categorize them as ‘Invalid’, and these ‘Invalid’ responses mostly include some city names instead of a country name (e,g., Rome, Tampa, Paris, etc.).
 Appendix C Extended Results
+Figure 7: Latin Personas’ RP average across all models
+Model $\tau$ Distance
+Max
+Distance
+Normalized
+Distance (%)
+GPT-4o 577.37 2775 20.80
+Llama3.1-70B 597.25 3240 18.43
+Mistral-7B 456.52 1711 26.68
+Gemma2-27B 400.51 1653 24.22
+Table 8: Kendall distance of general persona vs. all nation-specific personas together for all models. Max distance means perfect disagreement.
 Table 8 shows the Kendall $\tau$ distance between general persona and nationality-assigned personas.
-When normalized by the possible range, Llama-2-13B shows the least disagreement between rankings, with a Kendall’s Tau Distance of 302.32 out of a possible 2415 (approximately 12.5%).
-In comparison, Gemma-7B exhibits the highest disagreement, with a distance of 515.15 out of 1653 (approximately 31.2%).
-GPT-4o and Mistral-7B have distances of 279.58 (20.3%) and 169.51 (18.8%) out of their respective ranges.
-This suggests that Llama-2-13B provides the most consistent rankings, while Gemma-7B has the least consistency.
-[^9] The distance must be normalized because the raw Kendall $\tau$ distance is sensitive to the number of items compared and each model generated a different set of unique nations in the course of the experiment.
-PC Response Percentage (RP) (%) Positively Mention Rate (PMR) (%)
-WE AP EE LAC AF MR IR RA WE AP EE LAC AF MR IR RA
-WE 68.21 6.58 1.77 2.54 1.08 3.55 12.27 3.95 58.90 42.60 11.11 29.52 15.85 48.57 32.74 23.66
-AP 23.32 52.48 1.67 0.24 1.01 4.67 10.73 5.85 62.42 54.65 21.32 27.06 30.20 44.08 39.45 26.14
-EE 24.96 5.64 47.79 0.21 0.38 3.78 12.03 5.17 70.60 51.56 47.21 45.57 27.27 44.63 44.84 23.15
-LAC 26.04 5.29 0.69 48.18 0.46 4.50 11.50 3.29 61.38 67.91 9.93 49.65 41.49 48.14 41.05 19.21
-AF 21.43 6.71 0.63 0.82 47.58 7.12 10.80 4.86 66.59 68.78 11.43 40.39 50.37 42.79 39.94 21.80
-Table 7: Persona region results averaged across all models.
-Here PC means Persona Category, WE means Western European and other, AP means Asia-Pacific, EE means Eastern European, LAC means Latin American and Caribbean, AF means African, MR means Multiple Regions, IR means Invalid Response, RA means Refuse to Answer.
+When normalized by the possible range, Llama-3.1-70B shows the least disagreement between rankings, with a Kendall’s Tau Distance of 597.25 out of a possible 3240 (approximately 18.5%).
+In comparison, Mistral-7B exhibits the highest disagreement, with a distance of 456.52 out of 1711 (approximately 26.7%).
+GPT-4o and Gemma2-27B have distances of 577.37 (20.8%) and 400.51 (24.2%) out of their respective ranges.
+This suggests that Llama-3.1-70B provides the most consistent rankings, while Mistral-7B has the least consistency [^11] The distance must be normalized because the raw Kendall $\tau$ distance is sensitive to the number of items compared and each model generated a different set of unique nations in the course of the experiment..
 Appendix D Human Perception Vs Models Perception
-CP M GPT-4o Mistral Gemma Llama-2 Human
-Canadian 99.0 80.5 100 73.0 57.00
-Polish 100 70.5 91.7 74.5 93.00
-British 91.5 84.5 54.8 81.0 59.00
-Italian 96.5 70.0 79.7 67.7 60.00
-German 100 73.5 54.6 57.0 57.00
-Swedish 100 58.0 71.4 65.0 55.00
-Indian 100 75.5 69.7 57.4 65.00
-Japanese 100 68.5 62.5 53.3 73.00
-Hungarian 100 79.0 70.4 61.5 44.00
-French 99.0 71.0 100 55.4 52.00
-Mean $\Delta$ 37.1 17.0 17.7 12.2 -
-$\rho$ 0.09 -0.23 0.16 0.17 -
+CP M GPT-4o Mistral Gemma Llama Human
+Canadian 100.0 99.66 100.0 100.0 57.00
+Polish 100.0 100.0 100.0 100.0 93.00
+British 99.67 100.0 100.0 100.0 59.00
+Italian 100.0 99.65 100.0 100.0 60.00
+German 100.0 98.89 94.94 100.0 57.00
+Swedish 100.0 100.0 100.0 100.0 55.00
+Indian 100.0 100.0 100.0 100.0 65.00
+Japanese 100.0 100.0 100.0 100.0 73.00
+Hungarian 100.0 99.65 100.0 96.77 44.00
+French 100.0 99.28 78.95 64.24 52.00
+Mean $\Delta$ 38.467 38.213 35.889 34.601 -
+$\rho$0.058 0.579 0.412 0.685 -
 Table 9: Human perception Vs different models’ perception towards the United States after running the same experiment as the human experiment set-up.
@@
 Here, CP stands for Country Persona, M stands for Model.
-CN M GPT-4o Mistral Gemma Llama-2 Human
-Canada 100 51.5 100 92.0 88.00
-Russia 0.0 36.5 0.0 16.5 9.00
-UK 100 54.0 100 87.0 86.00
-Iran 0.0 26.5 8.0 2.5 15.00
-Iraq 1.0 37.0 4.5 0.0 17.00
-Mexico 100 47.0 100 94.5 59.00
-India 100 56.0 100 92.5 70.00
-Japan 100 53.0 100 97.5 81.00
-N. Korea 0.0 20.5 0.0 1.5 9.00
-France 100 53.5 100 96.5 83.00
-Mean $\Delta$ 18.2 22.3 17.1 13.8 -
-$\rho$ 0.88 0.78 0.87 0.59 -
+CN M GPT-4o Mistral Gemma Llama Human
+Canada 100.0 100.0 100.0 100.0 88.00
+Russia 0.0 3.0 0.0 0.0 9.00
+UK 100.0 100.0 100.0 100.0 86.00
+Iran 0.0 0.67 0.0 0.0 15.00
+Iraq 1.49 0.0 0.0 0.0 17.00
+Mexico 100.0 100.0 100.0 100.0 59.00
+India 100.0 100.0 100.0 100.0 70.00
+Japan 100.0 100.0 100.0 100.0 81.00
+N. Korea 0.0 0.0 0.0 0.0 9.00
+France 100.0 100.0 100.0 100.0 83.00
+Mean $\Delta$ 18.151 17.933 18.3 18.3 -
+$\rho$ 0.877 0.801 0.855 0.855 -
 Table 10: American Persona’s perception towards other countries after running the same experiment as the human experiment set-up.
@@
 Here, CN stands for Country Name, M stands for Model
-To get the results in Table 8, we ask the same questions to the LLMs as the Pew Research Center asked people to get human data.
-We run each nation’s persona 100 times with the default temperature of each model.
-Then we increase the temperature set it to 1.25, and run each nation’s persona 100 times.
-So, we have 200 data for each nation’s persona.
-See Table 10 for the question we used.
-Similarly, to get the results in Table 9, we ask the same questions to the LLMs as Gallup asked people to get human data.
-We also have 200 data for each country.
-See Table 10 for the exact question we used.
-Experimental set-up for Getting Results Shown in Table 8 and Table 9
-Table 8: Adopt the identity of a [demonyms] person.
-Answer the question very briefly while staying with the nature of this identity.
-What is your overall opinion of the United States?
-Is it favorable or unfavorable?
-You must answer either favorable or unfavorable or simply not answer.
-Table 9: Adopt the identity of an American person.
-Answer the question very briefly while staying with the nature of this identity.
-What is your overall opinion of [country name]?
-Is it very favorable, mostly favorable, mostly unfavorable or very unfavorable?.
-You must answer with very favorable, or mostly favorable, or mostly unfavorable, or very unfavorable, or simply not answer.
-Table 10: Questionaries used to get the results for Table 8 and Table 9. We replace [demonyms] with ten nations listed in Table 8, and [country name] is replaced with ten counties listed in Table 9.
-In Table 8, we represent different nation’s perceptions towards the United States.
+CP M
+Human Like
+Experiment’s
+Results for
+GPT-4o
+Our Primary
+Experiments
+Results for
+GPT-4o
+Human Like
+Experiment’s
+Results for
+Mistral
+Our Primary
+Experiments
+Results for
+Mistral
+Human Like
+Experiment’s
+Results for
+Gemma
+Our Primary
+Experiments
+Results for
+Gemma
+Human Like
+Experiment’s
+Results for
+Llama
+Our Primary
+Experiments
+Results for
+Llama
+Canadian 100.0 37.04 99.66 36.60 100.0 15.19 100.0 38.60
+Polish 100.0 83.33 100.0 57.20 100.0 60.00 100.0 95.52
+British 99.67 53.33 100.0 52.54 100.0 43.93 100.0 71.43
+Italian 100.0 46.15 99.65 48.42 100.0 43.33 100.0 71.13
+German 100.0 23.08 98.89 32.28 94.94 24.24 100.0 69.00
+Swedish 100.0 61.11 100.0 48.36 100.0 28.89 100.0 59.22
+Indian 100.0 66.67 100.0 59.52 100.0 48.78 100.0 68.09
+Japanese 100.0 60.00 100.0 32.14 100.0 32.35 100.0 37.86
+Hungarian 100.0 66.67 99.65 66.46 100.0 44.44 96.77 87.10
+French 100.0 19.61 99.28 41.25 78.95 29.69 64.24 67.59
+Table 11: Other nations’ perception towards the United States, comparing human-like experiments and our primary experiment’s results.
+All results are presented in PMR (%).
+Here, CP stands for Country Persona, M stands for Model.
+CN M
+Human Like
+Experiment’s
+Results for
+GPT-4o
+Our Primary
+Experiment’s
+Results for
+GPT-4o
+Human Like
+Experiment’s
+Results for
+Mistral
+Our Primary
+Experiment’s
+Results for
+Mistral
+Human Like
+Experiment’s
+Results for
+Gemma
+Our Primary
+Experiment’s
+Results for
+Gemma
+Human Like
+Experiment’s
+Results for
+Llama
+Our Primary
+Experiment’s
+Results for
+Llama
+Canada 100.0 100.0 100.0 80.0 100.0 98.55 100.0 84.85
+Russia 0.0 18.18 3.0 4.0 0.0 5.36 0.0 11.36
+UK 100.0 0.0 100.0 0.0 100.0 33.33 100.0 0.0
+Iran 0.0 0.0 0.67 0.0 0.0 0.0 0.0 0.0
+Iraq 1.49 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Mexico 100.0 0.0 100.0 20.0 100.0 0.0 100.0 5.0
+India 100.0 0.0 100.0 9.09 100.0 50.0 100.0 25.0
+Japan 100.0 92.86 100.0 85.71 100.0 89.80 100.0 85.29
+N. Korea 0.0 0.0 0.0 0.0 0.0 1.30 0.0 1.59
+France 100.0 46.15 100.0 52.08 100.0 27.78 100.0 44.58
+Table 12: American Persona’s perception towards other countries, comparing human-like experiments and our primary experiment’s results.
+All results are presented in PMR (%).
+Here, CN stands for Country Name, M stands for Model.
+In Table 9, we represent different nations’ perceptions towards the United States at the specific persona-model level.
 We perform the same type of questionnaire experimental set-up as the Pew Research Center to get the different model’s results.
-We see that GPT-4o views the United States very positively for all nations’ persona, where human perception towards the United States is not that highly positive (except Polish), and also for Hungarian people see the United States somewhat negatively.
+We see that GPT-4o treated the United States very positively for all nations’ persona, where human perception towards the United States is not that highly positive (except Polish), and also for Hungarian people see the United States somewhat negatively.
 For other models, we notice many variations of results.
-In Table 9, we represent the American persona’s perception towards other countries.
-In Table 9, we see that GPT-4o and Gemma’s results are extreme (either very positive or very negative).
+In Table 10, we represent the American persona’s perception towards other countries at the specific country-model level.
+In Table 10, we see that all models’ results are extreme (either very positive or very negative).
 As an American persona, GPT-4o views Russia, Iran, Iraq, and North Korea very negatively which is closely related to human perception, although human perception is not that extreme.
-Llama-2’s results are more closely related to human perception.